# FedFDP: Federated Learning with Fairness and Differential Privacy

### v1

@@ 4 Proposed Algorithm: FedFair
  • Local (lines 1-1): For $E$ epochs and each batch $b$, we compute the adaptive learning rate $\eta_{i}$ in lines 1-1. For line 1, we execute gradient descent by using $\eta_{i}$.
@@
Furthermore, as the iterations advance, the model progressively prioritizes fairness, leading to a reduction in $\eta_{i}$. In comparison to training methods with a fixed learning rate, learning rate decay can accelerate the convergence of gradient-based methods Zinkevich (2003).
Figure 2: FedFDP framework
5 Proposed Algorithm: FedFDP
@@
In this section, we propose the FedFDP by integrating differential privacy protection base on FedFair. The flow of FedFDP algorithm as shown in Fig. 2. The FedFDP algorithm adds the following two additional processes to FedFair. Firstly, in Section 5.1, we present a fair-clipping strategy that not only achieves sensitivity but also allows for adjustments to the level of fairness. Furthermore, in Section 5.2, we expand the use of differential privacy protection to the additional loss values that local clients need to upload, suggesting an adaptive clipping method to maximize utility. The overall FedFDP algorithm is described in Section 5.3.
5.1 Fair-clipping Strategy for Gradient
In the general DPSGD algorithm Fu et al. (2023); Abadi et al. (2016); Bu et al. (2022); Yang et al. (2022), per-sample clipping is an essential process for obtaining sensitivity. In order to be compatible with per-sample clipping, we expand our objective function of client $i$ in FedFair, Equation (7), into Equation (10):
$\min_{\mathbf{w}_{t}^{i}}H_{i}(\mathbf{w}_{t}^{i},\xi_{j})\triangleq F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\frac{\lambda}{2}\left(F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t})\right)^{2},$ (10)
Next, we compute the gradient of Equation (10):
$\nabla H_{i}(\mathbf{w}_{t}^{i},\xi_{j})$ $=\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\lambda\cdot\Delta_{loss}^{j}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$
$=\left(1+\lambda\cdot\Delta_{loss}^{j}\right)\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j}),$ (11)
where $\Delta_{loss}^{j}=F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t})$.
To achieve differential privacy protection, we replace loss function gradient by Equation (5.1) based on Equations (3)-(5), and reorganize the gradient descent equation for client $i$:
$\mathbf{w}_{t+1}^{i}$ $=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}\frac{\left(1+\lambda\cdot\Delta^{j}_{loss}\right)\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})}{\max\left(1,\frac{\left(1+\lambda\cdot\Delta^{j}_{loss}\right)\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}{C}\right)}$
$\quad+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]$
$=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})],$ (12)
where:
$C^{i,j}_{t}=\min\left(1+\lambda\cdot\Delta^{j}_{loss},\frac{C}{\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}\right).$ (13)
Hence, our strategy can be regarded as an fair-clipping technique, tailored to meet the demands of differential privacy while also nurturing fairness within the model. When $\lambda=0$, Equation (13) corresponds to the traditional DP clipping method, and FedFDP no longer ensures fairness. For large $C$, Equation (13) simplifies to the first term, prompting FedFDP to adjust towards enhanced fairness. Conversely, for small $C$, Equation (13) simplifies to the second term, resulting in the gradient being clipped to the bound $C$. Our conclusions can be verified from the experiments 5.
5.2 Adaptive Clipping Method for Loss
Since the calculation of $\Delta_{loss}^{j}$ requires clients to upload $F_{i}(\mathbf{w}_{t})$, we need to incorporate differential privacy to ensure that the algorithm does not compromise privacy. As shown in lines 2-2 of Algorithm 2, We need to clip $F_{i}(\mathbf{w}_{t})$ to the interval $[0,C_{l}^{i,t}]$, and then add Gaussian noise with a mean of 00 and a standard deviation of $C_{l}^{i,t}\cdot\sigma_{l}$ to ensure differential privacy. However, setting an appropriate clipping bound $C_{l}^{i,t}$ for $F_{i}(\mathbf{w}_{t})$ is challenging. A large $C_{l}^{i,t}$ helps avoid loss of fidelity to the original values, but it also leads to large injected noise.
Based on the data heterogeneity of federated learning, we propose an adaptive clipping method for $F_{i}(\mathbf{w}_{t})$. For the $i$-th client, the adaptive clipping method uses the $differentially$ $private$ $mean$ $of$ $the$ $previous$ $round$ as the clipping bound for the current round. The clipping bound $C_{l}^{i,t}$ for round $t$ and client $i$ is defined as follows, when $F_{i}(\mathbf{w}_{t-1})$ is the individual loss of the previous round and $\sigma^{t-1}_{l}$ is the individual noise multiplier of the previous round.
$C_{l}^{i,t}=\frac{\sum_{j=1}^{|\mathcal{B}_{i}|}\operatorname{clip}\left(F_{i}(\mathbf{w}_{t-1},\xi_{j})\right)+\mathcal{N}\left(0,({C_{l}^{i,t-1}}\cdot{\sigma_{l}})^{2}\right)}{|\mathcal{B}_{i}|},$ (14)
where
$\operatorname{clip}\left(F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j})\right)=\min({C_{l}^{i,t-1}},\max(0,F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j}))),$ (15)
5.3 Overall Algorithm
We utilize the equation form presented in (12) to enhance Algorithm 2. Compared with the Algorithm 1, FedFDP solely necessitates the application of differential privacy to gradient and loss.
@@ 5.3 Overall Algorithm
Input: loss function $F(\mathbf{w})$. Parameters: learning rate $\eta$, noise multiplier for training $\sigma$, noise multiplier for loss $\sigma_{l}$, fairness hyperparameter $\lambda$, batch sample ratio $q$, original clipping bound $C$.
@@
1 Initialize $\mathbf{w}_{0},F(\mathbf{w}_{0}),C_{l}^{0}=\text{Initial()}$();
2SERVER
@@ 14       Sample randomly a batch $\mathcal{B}_{i}$ with probability $q$;
16             $\Delta_{loss}^{j}=F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t})$;
@@ 18             $\tilde{g}_{t}^{i,j}=C_{t}^{i,j}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$;
20      $\mathbf{w}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}\tilde{g}_{t}^{i,j}+\sigma C\mathcal{N}(0,\mathbf{I})\right)$;
@@ 24             $f_{i,j}=\min({C_{l}^{i,t}},\max(0,F_{i}(\mathbf{w}_{t+1}^{i},\xi_{j})))$;
25
26      $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})\leftarrow\frac{1}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}f_{i,j}+\sigma_{l}C_{l}^{i,t}\mathcal{N}(0,\mathbf{I})\right)$;
@@
6 Utility Analysis
In this section, we examine the convergence of the FedFDP algorithm to identify an optimal $\lambda$ that minimizes the convergence bound.
6.1 Convergence Analysis
@@
Let $\xi_{j}$ be sampled from the $i$-th device’s local data uniformly at random. The expected squared norm of stochastic gradients is uniformly bounded, i.e., $\left\|\nabla F_{i}\left(\mathbf{w}_{t}^{i},\xi_{j}\right)\right\|\leq G$ for all $i=1,\cdots,N$ , $t=1,\cdots,T-1$ and $j=1,\cdots,|\mathcal{B}_{i}|$
@@
$\mathbb{E}[F(\mathbf{w}_{t})]-F^{*}\leq\frac{L}{2t}\left(\frac{A}{\mu^{2}(2C_{t}-1)}+\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}\right),$ (16)
@@
The meaning of $\Gamma$ is consistent with Li et al. (2020c), where a larger value of $\Gamma$ indicates that the data among different clients is more Non-IID.
@@
proof sketch: We investigate the relationship between $\mathbb{E}\|\mathbf{w}_{t+1}-\mathbf{w}^{*}\|^{2}$ and $\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$, and then use mathematical induction to obtain an upper bound for $\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$. Finally, by utilizing Assumption 1, we derive (16). For the detailed proof, please refer to Appendix A.1.
6.2 Optimal $\lambda$
@@
When $t\rightarrow T$ and $T$ is sufficiently large, $\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|\rightarrow 0$, $C_{t}^{i,j}=1+\lambda\Delta_{loss}^{j}$. We assume $Q_{0}\leq\Delta_{loss}^{j}\leq Q_{1}$, then $1+Q_{0}\lambda\leq C_{t}^{i,j}\leq 1+Q_{1}\lambda$. We rearrange the convergence upper bound in (16) as a function of $\lambda$,
$P(\lambda)=\frac{L}{2\mu T}\frac{a_{1}\lambda^{3}+a_{2}\lambda^{2}+a_{3}\lambda+a_{4}}{2Q_{0}\lambda+1},$ (17)
where $a_{1}=G^{2}Q_{1}^{3}$, $a_{2}=6G^{2}Q_{1}^{2}$, $a_{3}=9Q_{1}G^{2}+2L\Gamma Q_{1}+2Q_{0}\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}$, $a_{4}=4G^{2}+2L\Gamma+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}+\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}$.
We sketched a rough graph of $P(\lambda)$, where $\lambda\in[0,+\infty)$. Upon observing Fig. 3, it becomes evident that when $a_{1}$,$a_{2}$,$a_{3}$,$a_{4}$ greater than 00, there exists an optimal value of $\lambda$ that minimizes the right-hand side of Equation (16), aligning with the experimental results presented in Fig. 4.
Figure 3: An example of $P(\lambda)$
7 Privacy Analysis
Because each client’s locally trained pair of data samples has the same sampling rate and noise multiplier in each round, each client has the same privacy loss in each round. As a result, we only assess each client’s privacy loss from the perspective of a single client. Following that, we examine the privacy loss of FedPDF for the $i$-th client as follows.

### v2

@@ 4. Proposed Algorithm: FedFair
  • Local (lines 1-1): For each batch $b$, we compute the adaptive learning rate $\eta_{i}$ in lines 1-1. For line 1, we execute gradient descent by using $\eta_{i}$.
@@
Furthermore, as the training process advances, the model progressively prioritizes fairness, leading to a reduction in $\eta_{i}$. In comparison to training methods with a fixed learning rate, learning rate decay can accelerate the convergence of gradient-based methods (Zinkevich, 2003).
5. Proposed Algorithm: FedFDP
@@
In this section, we propose the FedFDP by integrating differential privacy protection base on FedFair. The flow of FedFDP algorithm as shown in Figure 2. The FedFDP algorithm adds the following two additional processes to FedFair. Firstly, in Section 5.1, we present a fair-clipping strategy that not only achieves sensitivity but also allows adjustments to the level of fairness. Furthermore, in Section 5.2, we expand the use of differential privacy protection to the additional loss values that local clients need to upload, suggesting an adaptive clipping method to maximize utility. The overall FedFDP algorithm is described in Section 5.3.
@@
Input: loss function $F(\mathbf{w})$. Parameters: learning rate $\eta$, noise multiplier for gradient $\sigma$, noise multiplier for loss $\sigma_{l}$, fairness hyperparameter $\lambda$, batch sample ratio $q$, original clipping bound $C$.
@@
1 Initialize $\mathbf{w}_{0},F(\mathbf{w}_{0}),C_{l}^{0}$;
2 SERVER
@@ 14       Sample randomly a batch $\mathcal{B}_{i}$ with probability $q$;
16             $\Delta_{i}^{j}=F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t})$;
@@ 18             $\tilde{g}_{t}^{i,j}=C_{t}^{i,j}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$;
20      $\mathbf{w}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}\tilde{g}_{t}^{i,j}+\sigma C\cdot\mathcal{N}(0,\mathbf{I})\right)$;
21
@@ 26            $f_{i,j}=\min({C_{l}^{i,t}},\max(0,F_{i}(\mathbf{w}_{t+1}^{i},\xi_{j})))$;
27
28      $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})\leftarrow\frac{1}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}f_{i,j}+\sigma_{l}C_{l}^{i,t}\cdot\mathcal{N}(0,\mathbf{I})\right)$;
@@
5.1. Fair-clipping Strategy for Gradient
In the general DPSGD algorithm (Abadi et al., 2016; Fu et al., 2023; Bu et al., 2022; Yang et al., 2022), per-sample clipping is an essential process for obtaining sensitivity. In order to be compatible with per-sample clipping, initially, we expand our objective function of client $i$ in FedFair, Equation (9), into Equation (12):
(12) $\min_{\mathbf{w}_{t}^{i}}H_{i}(\mathbf{w}_{t}^{i},\xi_{j})\triangleq F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\frac{\lambda}{2}\left(F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t}^{i})\right)^{2},$
Next, we compute the gradient of Equation (12):
$\nabla H_{i}(\mathbf{w}_{t}^{i},\xi_{j})$ $=\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\lambda\cdot\Delta_{i}^{j}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$
(13) $=\left(1+\lambda\cdot\Delta_{i}^{j}\right)\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j}),$
where $\Delta_{i}^{j}=F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t}^{i})$.
To achieve differential privacy protection, we replace the gradient of loss function by Equation (5.1) based on Equations (4)-(6), and reorganize the gradient descent equation for client $i$:
$\mathbf{w}_{t+1}^{i}$ $=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}\frac{\left(1+\lambda\cdot\Delta_{i}^{j}\right)\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})}{\max\left(1,\frac{\left(1+\lambda\cdot\Delta_{i}^{j}\right)\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}{C}\right)}$
$\quad+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]$
(14) $=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})],$
where:
(15) $C^{i,j}_{t}=\min\left(1+\lambda\cdot\Delta_{i}^{j},\frac{C}{\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}\right).$
Hence, our strategy can be regarded as an fair-clipping technique, tailored to meet the demands of differential privacy while also nurturing fairness within the model. When $\lambda=0$, Equation (15) corresponds to the traditional DP clipping method, and FedFDP no longer ensures fairness. For large $C$, Equation (15) simplifies to the first term, prompting FedFDP to adjust towards enhanced fairness. Conversely, for small $C$, Equation (15) simplifies to the second term, resulting in the gradient being clipped to the bound $C$. Our conclusions can be verified from the experiments (Figure 4).
5.2. Adaptive Clipping Method for Loss
Since the calculation of $\Delta_{i}^{j}$ requires clients to upload $F_{i}(\mathbf{w}_{t}^{i})$, we need to incorporate differential privacy to ensure that the algorithm does not compromise privacy. As shown in lines 2-2 of Algorithm 2, We need to clip $F_{i}(\mathbf{w}_{t}^{i})$ to the interval $[0,C_{l}^{i,t}]$, and then add Gaussian noise with a mean of 00 and a standard deviation of $\sigma_{l}\cdot C_{l}^{i,t}$ to ensure differential privacy. However, setting an appropriate clipping bound $C_{l}^{i,t}$ for $F_{i}(\mathbf{w}_{t}^{i})$ is challenging. A large $C_{l}^{i,t}$ avoids loss of fidelity to the original values, but it also leads to large injected noise.
Based on the data heterogeneity of federated learning, we propose an adaptive clipping method for $F_{i}(\mathbf{w}_{t}^{i})$. For the $i$-th client, the adaptive clipping method uses the $differentially$ $private$ $mean$ $of$ $the$ $previous$ $round$ as the clipping bound for the current round. The clipping bound $C_{l}^{i,t}$ for round $t$ and client $i$ is defined as follows, when $F_{i}(\mathbf{w}_{t-1}^{i})$ is the individual loss of the previous round and $\sigma^{t-1}_{l}$ is the individual noise multiplier of the previous round.
(16) $C_{l}^{i,t}=\frac{\sum_{j=1}^{|\mathcal{B}_{i}|}\operatorname{clip}\left(F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j})\right)+\mathcal{N}\left(0,({C_{l}^{i,t-1}}\cdot{\sigma_{l}})^{2}\right)}{|\mathcal{B}_{i}|},$
where
(17) $\operatorname{clip}\left(F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j})\right)=\min({C_{l}^{i,t-1}},\max(0,F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j}))).$
5.3. Overall Algorithm
We utilize the equation form Equation (14) to enhance Algorithm 2. Compared with the Algorithm 1, FedFDP solely necessitates the application of differential privacy to gradient and loss.
@@ 5.3. Overall Algorithm
6. Optimal Fairness Parameter
In this section, we examine the convergence of the FedFDP algorithm to identify an optimal fairness parameter $\lambda$ that minimizes the convergence bound.
6.1. Convergence Analysis
@@
Let $\xi_{j}$ be sampled from the $i$-th device’s local data uniformly at random. The expected squared norm of stochastic gradients is uniformly bounded, i.e., $\left\|\nabla F_{i}\left(\mathbf{w}_{t}^{i},\xi_{j}\right)\right\|\leq G$ for all $i=[1,\cdots,N]$ , $t=[1,\cdots,T-1]$ and $j=[1,\cdots,|\mathcal{B}_{i}|]$
@@
(18) $\mathbb{E}[F(\mathbf{w}_{t})]-F^{*}\leq\frac{L}{2t}\left(\frac{A}{\mu^{2}(2C_{t}-1)}+\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}\right),$
@@
The meaning of $\Gamma$ is consistent with (Li et al., 2020a), where a larger value of $\Gamma$ indicates that the data among different clients is more Non-IID.
@@
proof sketch: We investigate the relationship between $\mathbb{E}\|\mathbf{w}_{t+1}-\mathbf{w}^{*}\|^{2}$ and $\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$, and then use mathematical induction to obtain an upper bound for $\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$. Finally, by utilizing Assumption 1, we derive (18). For the detailed proof, please refer to Appendix A.1333Due to the lengthy proof, please refer to https://anonymous.4open.science/r/FedFDP-5607/APPENDIX/FedFDP_appendix.pdf for the appendix..
6.2. Optimal $\lambda$
Theorem 2.
There exists a computable optimal $\lambda$, denoted as $\lambda^{*}$, which ensures the fastest convergence of the Algorithm 2.
@@ 6.2. Optimal $\lambda$
Proof. When $t\rightarrow T$ and $T$ is sufficiently large, $\|\nabla F_{i}(\mathbf{w}_{t}^{i},x_{j})\|\rightarrow 0$, $C_{t}^{i,j}=1+\lambda\Delta_{i}^{j}$. Assuming $Q_{0}\leq\Delta_{i}^{j}\leq Q_{1}$, then $1+Q_{0}\lambda\leq C_{t}^{i,j}\leq 1+Q_{1}\lambda$. We rearrange the convergence upper bound in Equation (18) to obtain a function in terms of $\lambda$,
(19) $P(\lambda)=\frac{L}{2\mu T}\frac{a_{1}\lambda^{3}+a_{2}\lambda^{2}+a_{3}\lambda+a_{4}}{a_{5}\lambda+1},$
where $a_{1}=G^{2}Q_{1}^{3}$, $a_{2}=6G^{2}Q_{1}^{2}$, $a_{3}=9Q_{1}G^{2}+2L\Gamma Q_{1}+2Q_{0}\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}$, $a_{4}=4G^{2}+2L\Gamma+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}+\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}$, $a_{5}=2Q_{0}$.
Let $\mathcal{F}(\lambda)=\frac{a_{1}\lambda^{3}+a_{2}\lambda^{2}+a_{3}\lambda+a_{4}}{a_{5}\lambda+1}$, then $P(\lambda)=\frac{L}{2\mu T}\mathcal{F}(\lambda)$, where $\frac{L}{2\mu T}$ can be considered a constant term. Obviously, when $\mathcal{F}(\lambda)$ is minimized, $P(\lambda)$ is also minimized. The derivative of $\mathcal{F}(\lambda)$ is as follows:
$\mathcal{F}^{\prime}(\lambda)$ $=\frac{(3a_{1}\lambda^{2}+2a_{2}\lambda+a_{3})(a_{5}\lambda+1)}{(a_{5}\lambda+1)^{2}}-\frac{a_{5}(a_{1}\lambda^{3}+a_{2}\lambda^{2}+a_{3}\lambda+a_{4})}{(a_{5}\lambda+1)^{2}}$
(20) $=\frac{2a_{1}a_{5}\lambda^{3}+(a_{2}a_{5}+3a_{1})\lambda^{2}+2a_{2}\lambda+a_{3}-a_{4}a_{5}}{(a_{5}\lambda+1)^{2}}.$
Let the denominator of $\mathcal{F}^{\prime}(\lambda)$ be a new function:
(21) $\mathcal{G}(\lambda)=2a_{1}a_{5}\lambda^{3}+(3a_{1}+a_{2}a_{5})\lambda^{2}+2a_{2}\lambda+a_{3}-a_{4}a_{5}.$
Since $(a_{5}\lambda+1)^{2}>0$, the sign of $\mathcal{F}^{\prime}(\lambda)$ is consistent with that of $\mathcal{G}(\lambda)$.
Using the discriminant of a cubic function $\Delta=b^{2}-3ac$ (a general cubic function is expressed as $a\lambda^{3}+b\lambda^{2}+c\lambda+d$),
$\Delta$ $=(3a_{1}+a_{2}a_{5})^{2}-3(2a_{1}a_{5}\cdot 2a_{2})$
$=9a_{1}^{2}+a_{2}^{2}a_{5}^{2}+6a_{1}a_{2}a_{5}-12a_{1}a_{2}a_{5}$
(22) $=(3a_{1}-a_{2}a_{5})^{2}\geq 0.$
The derivative of $\mathcal{G}(\lambda)$ is given by
(23) $\mathcal{G}^{\prime}(\lambda)=6a_{1}a_{5}\lambda^{2}+2(a_{2}a_{5}+3a_{1})\lambda+2a_{2}.$
It is evident that when $\lambda>0$, $\mathcal{G}^{\prime}(\lambda)>0$.
If $\mathcal{F}(\lambda)$ has an extremum, then $\mathcal{F}^{\prime}(\lambda)$ has a root, meaning $\mathcal{G}(\lambda)$ has a root. Substituting $\lambda=0$, $\mathcal{G}(0)=a_{3}-a_{4}a_{5}$, evidently when $a_{3}<a_{4}a_{5}$, $\mathcal{G}(\lambda)$ has a root.
Let $\lambda=x-\frac{a_{2}a_{5}+3a_{1}}{6a_{1}a_{5}}$, replacing $\lambda$ with $x$ in $\mathcal{G}(\lambda)$ yields
(24) $\mathcal{G}(x)=x^{3}+a_{6}x+a_{7},$
where
  • $a_{6}=\frac{-(3a_{1}-a_{2}a_{5})^{2}}{12a_{1}^{2}a_{5}^{2}}$,
  • $a_{7}=\frac{108a_{1}^{2}a_{5}^{2}(a_{3}-a_{4}a_{5})-36(a_{1}a_{2}a_{5})(3a_{1}+a_{2}a_{5})+2(3a_{1}+a_{2}a_{5})^{3}}{216a_{1}^{3}a_{5}^{3}}$.
Using the root-finding formula to calculate $\mathcal{G}(x)=0$ gives the solutions $x_{1},x_{2},x_{3}$.
[…]
where imaginary number $\omega=\frac{-1+\sqrt{3}\mathbf{i}}{2}$, $\lambda_{1,2,3}=x_{1,2,3}-\frac{a_{2}a_{5}+3a_{1}}{6a_{1}a_{5}}$.
Furthermore, since $\lambda_{1}\lambda_{2}\lambda_{3}=-\frac{a_{3}-a_{4}a_{5}}{2a_{1}a_{5}}>0$, $\lambda$ has only one positive real root. If the equation has $k$ real roots, the optimal solution for $\lambda$ is $\lambda^{*}=\max_{k}(\lambda_{k})$.
In conclusion, $\mathcal{F}^{\prime}(\lambda)\geq 0\iff G(\lambda)\geq 0\iff\lambda\geq\lambda^{*}$, therefore $\mathcal{F}(\lambda)$ has a minimum value $\mathcal{F}(\lambda^{*})$, and $P(\lambda)$ has a minimum value $P(\lambda^{*})$.
7. Privacy Analysis
7.1. Privacy Loss of $\mathbf{w}_{t+1}^{i}$
Theorem 1.
(25) $R_{model}^{i}(\alpha)=\frac{T}{\alpha-1}\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-i}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma^{2}}\right),$
where $\sigma$ is noise multiplier of the $\mathbf{w}_{t+1}^{i}$, and $\alpha>1$ is the order.
The detail of proof please refer Appendix A.2.
7.2. Privacy Loss of $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$
Theorem 2.
(26) $R_{loss}^{i}(\alpha)=\frac{T}{\alpha-1}\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-i}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma_{l}^{2}}\right),$
The proof is similar to that of the $\mathbf{w}_{t+1}^{i}$, so we omit it.
7.3. Privacy Loss of FedFDP
Since both $\mathbf{w}_{T}^{i}$ and $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ access the training set, we need to combine their RDP sequentially using Definition 3, and then use Lemma 4 to convert it to $(\epsilon,\delta)$-DP. Lastly, we can get the Privacy loss of FedFDP as follows.